# Towards Scene Graph Anticipation

### v1

@@ 5.1 Object Representation Processing Unit
To achieve temporally consistent object representations across observed frames, our approach first extracts visual features using a pre-trained object detector and then further refines these features by processing them through an encoder. Specifically, We utilize a pre-trained Faster R-CNN[31] for extracting visual features ($\{\mathbf{v}^{t}_{i}\}_{i=1}^{N(t)}$), bounding boxes ($\{\mathbf{b}^{t}_{i}\}_{i=1}^{N(t)}$), and object category distributions ($\{\hat{\mathbf{c}}^{t}_{i}\}_{i=1}^{N(t)}$), for object proposals $\{o^{t}_{i}\}_{i=1}^{N(t)}$ in the observed frames. We then compute the matrix $\mathbf{V}_{i}$ by stacking visual features $\{\mathbf{v}_{i}^{t}\}_{t=1}^{T}$. We employ a transformer encoder [35] to aggregate temporal information across visual features from all frames. The input for our encoder consists of sequences $\mathbf{V}_{i}$, which serve simultaneously as queries $\mathbb{Q}$, keys $\mathbb{K}$ and values $\mathbb{V}$. Formally, the transformation process at the $n$-th encoder layer is given by
@@ 5.2 Spatial Context Processing Unit
To learn spatial context-aware relationship representations. We first construct relationship representations between interacting objects in a scene and further process them by passing through an encoder. Specifically, following [7], we construct relationship representation $\mathbf{z}_{ij}^{t}$ by concatenating visual and semantic features of the objects as follows:
@@
Here, $\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{W}_{3}$ represent the learnable weights of linear layers. $\mathbf{U}_{ij}^{t}$ represent processed feature maps of the union box computed by RoIAlign [31]. $\mathbf{S}_{i}^{t}$ and $\mathbf{S}_{j}^{t}$ correspond to the semantic embedding vectors of the object categories. Subsequently, we employ a transformer encoder to integrate spatial contextual information. Here, for each observed frame $\mathcal{I}^{t}$, we construct $\mathbf{Z}^{t}$ by stacking all relationship features $\{\mathbf{z}_{ij}^{t}\}_{ij}$ corresponding to objects observed in a frame. We then feed it as input for a transformer encoder which operates on queries $\mathbb{Q}$, keys $\mathbb{K}$, and values $\mathbb{V}$ that are derived from the same source, the preceding layer’s output or the initial relationship features. Formally, the transformation process at the $n$-th encoder layer is given by:
$\mathbf{Z}^{t}_{(n)}=\operatorname{SpatialEncoder}\left(\mathbb{Q}=\mathbb{K}=\mathbb{V}=\mathbf{Z}^{t}_{(n-1)}\right)$ (5)
@@ 5.3 Latent Dynamics Processing Unit
To understand spatio-temporal dynamics of the evolution of relationships between interacting objects, departing from traditional approaches that architectural variants of transformers [11, 34], we take a continuous time approach and learn a governing differential equation in latent space of relationship representations. Our LDPU contains two components: (1) Temporal Context Encoder and (2) Latent Dynamics Model.
5.3.1 Temporal Context Encoder.
To learn spatio-temporal context-aware relationship representations, we further process the output of SCPU by passing it through a transformer encoder that fine-tunes the representation. Specifically, we compute the matrix $\mathbf{Z}_{ij}$ by vertically stacking the relationship representations $\{\mathbf{z}_{ij}^{t}\}_{t=1}^{T}$. As described in Eq. 6, we feed $\mathbf{Z}_{ij}$ through an encoder that aggregates temporal information across all observed relationship representations. Here, the input $\mathbf{Z}_{ij}$, serves simultaneously as queries $\mathbb{Q}$, keys $\mathbb{K}$, and values $\mathbb{V}$.
@@
5.3.2 Latent Dynamics Models.
We begin by abstracting the complexities introduced by external factors that add uncertainty to the evolution of relationships between interacting objects and construct a model based on the premise that the core dynamics driving the changes in these relationships are governed by a non-linear deterministic process. Specifically, we leverage the expressive nature of NeuralODEs [3] in learning time-invariant vector fields from data that approximate the underlying non-linear deterministic process and propose SceneSayerODE. We cast the problem of anticipating relationship representations of interacting objects as an instance of the ODE-IVP, where the initial condition is given by the relationship representation of the last observed interaction of the object pair $\left(o_{i}^{t},o_{j}^{t}\right)$. The evolution of this relationship over time is then mathematically described as:
@@
In video data that is captured from a single viewpoint, we lose information due to issues such as blurry imagery and occlusions, leading to uncertain interpretations of the scenes. Therefore, it is crucial to integrate uncertainty into the modelling frameworks to represent these stochastic dynamics accurately. We assume the presence of a non-linear stochastic differential equation that governs this evolution and propose SceneSayerSDE which uses NeuralSDEs [17] to learn it from data. Specifically, we formulate the problem of anticipating future interactions as an SDE-IVP. Here, the initial conditions are set using the representations of the last observed representation object pair, thus facilitating a data-driven learning process that accounts for the inherent uncertainties. The evolution of this relationship over time is then described as:
@@ 5.4 Decoders
Predicate Classification Head. We employ two predicate classification heads in our approach; we use one for classifying relationship representations between objects of the observed scenes and the other for classifying anticipated relationship representations of the future scenes. Our predicate classification head is implemented using a two layer neural network. Bounding Box Regression Head. Our regression head takes the anticipated relationship representations as input and outputs bounding box information corresponding to the interacting objects. It is implemented using two layers of neural network. Object Classification Head. We integrate processed object representations obtained from the object encoder with object category distributions from the object detection process and pass them through a two-layer neural network to output refined object category distributions.
@@
Object Classification Loss. The classification effectiveness is evaluated using the standard cross-entropy loss, denoted as $\mathcal{L}_{(o)}^{t}$
$\mathcal{L}_{(o)}^{t}=-\sum_{c=1}^{|\mathcal{C}|}y_{o,c}^{t}\log(p_{o,c}^{t})$ (9)
Predicate Classification Loss. The model employs a dual-component loss system to understand and predict object relationships over time given as follows:
$\mathcal{L}_{\text{gen}}=\sum_{t=1}^{T}\sum_{ij}\mathcal{L}_{p^{t}_{ij}},\quad\mathcal{L}_{\text{ant}}=\sum_{t=T+1}^{T+H}\sum_{ij}\mathcal{L}_{p^{t}_{ij}},$ (10)
The first component, the Generation Loss focuses on classifying the relationships between pairs of objects $\left(o_{i}^{t},o_{j}^{t}\right)$ across all observed frames ($t\in[1,T]$). The second component, the Anticipation Loss $\mathcal{L}_{\text{ant}}$ is designed for future frames ($t\in[T+1,T+H]$) and applied on extrapolated relationship representations. Here $\mathcal{L}_{p^{t}_{ij}}$ is given by
$\mathcal{L}_{p^{t}_{ij}}(\hat{\mathbf{p}}^{t}_{ij})=\sum_{i\in\mathcal{P}^{+}}\sum_{j\in\mathcal{P}^{-}}\max(0,1-\hat{\mathbf{p}}^{t}_{ij}[j]+\hat{\mathbf{p}}^{t}_{ij}[i])$ (11)
Bounding Box Regression Loss. In our model, each anticipated relationship representation is input into a dedicated linear layer to calculate the predicted bounding boxes for objects. The deviation between these predicted bounding boxes and the ground truth bounding boxes is quantified using the Smoothed L1 Loss, denoted as $\mathcal{L}_{\text{boxes}}$
$\mathcal{L}_{\text{boxes}}^{t}=\sum_{i\in\text{boxes}}\text{L}_{smooth}(y_{i}^{t}-\hat{y}_{i}^{t})$ (12)
Reconstruction Loss. This loss function ensures that the reconstructed relationship representations closely mirror the outputs of temporal encoder in LDPU.
$\mathcal{L}_{recon}^{t}=\frac{1}{N(t)\times N(t)}\sum_{ij}^{(N(t)\times N(t))}\text{L}_{smooth}(\mathbf{z}_{ij}^{t}-\hat{\mathbf{z}}_{ij}^{t})$ (13)
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{gen}}+\lambda_{2}\mathcal{L}_{\text{ant}}+\lambda_{3}\sum_{t=T+1}^{T+H}\mathcal{L}_{\text{boxes}}^{t}+\lambda_{4}\sum_{t=1}^{T}\mathcal{L}_{(o)}^{t}+\lambda_{5}\sum_{t=T+1}^{T+H}\mathcal{L}_{recon}^{t}$ (14)
@@ 6 Experiments
Dataset. We apply the proposed method to anticipate future interactions on the Action Genome [15]. We pre-process the data and filter out videos with less than 3 annotated frames. Thus, we obtained 11.4K videos in total; we adhered to the train and test split provided by the dataset. The dataset encompasses 35 object classes and 25 relationship classes. These 25 relationship classes are grouped into three categories, namely: (1) Attention Relations comprise relationship classes which primarily describe attention of the subject towards the object, (2) Spatial Relations comprise relationship classes that describe the spatial relationship between two objects, and (3) Contacting Relations comprises relationship classes that indicate different ways the object is contacted.
@@
Remark. This dataset primarily includes videos featuring a single actor engaging with objects [^4] Adopting an actor-centric approach emphasises the actor’s importance in the scene and also aligns with the literature on VidSGG. While SceneSayer can anticipate relationship representations between any two objects in a scene, we specifically tailor it to address the dynamics between the actor (subject) and the objects with which they interact. in various real-world environments. In the context of the action genome, a subject-object pair can demonstrate multiple spatial and contacting relationships.
@@
Evaluation Metric. We evaluate our models using the standard Recall@K and meanRecall@K metrics, where $K$ is set to values within the set {10, 20, 50}. Recall@K metric assists in assessing the ability of our model to anticipate the relationships between observed objects in future frames. The long-tailed distribution of relationships in the training set [27] can generate biased scene graphs. While the performance on more common relationships can dominate the Recall@K metrics, the mean recall metric introduced in [4] is a more balanced metric that scores the model’s generalisation to all predicate classes. In the following sections, we will assess the performance of models by modifying the initial fraction of the video, denoted as $\mathcal{F}$ provided as input. We set $\mathcal{F}$ to 0.3, 0.5, 0.7, and 0.9 to facilitate a comprehensive understanding of the model’s proficiency in short-term and long-term relationship anticipation.
@@
Settings. We introduce three settings for evaluating models designed for the SGA task. Each setting is defined by the varying degrees of scene information made available to the model as input: (a) Action Genome Scenes (AGS): Here, the model’s input is limited to frames of the video. With a basic level of information, the model’s ability to interpret the scenes with minimal context is challenged. (b) Partially Grounded Action Genome Scenes (PGAGS): In this intermediate setting, the model receives additional information beyond the basic video frames. Specifically, it is supplied with precise bounding box details for active interacting objects within the scene. (c) Grounded Action Genome Scenes (GAGS): This setting provides the most comprehensive level of scene information. The model is given precise bounding box information and the categories of the active objects involved in the interaction. In video data, frames frequently encounter challenges such as object occlusions and blurry imagery, which complicate object detection and interpretation of interaction. In a setting where complete information is available, noise induced by these challenges significantly decreases. Consequently, we aim to evaluate the model’s ability to understand spatio-temporal dynamics.

### v2

@@ 5.1 Object Representation Processing Unit
To achieve temporally consistent object representations across observed frames, our approach first extracts visual features using a pre-trained object detector and then further enhances these extracted features by passing them through an encoder. Specifically, We utilize a pre-trained Faster R-CNN[ren2016faster] for extracting visual features ($\{\mathbf{v}^{t}_{i}\}_{i=1}^{N(t)}$), bounding boxes ($\{\mathbf{b}^{t}_{i}\}_{i=1}^{N(t)}$), and object category distributions ($\{\hat{\mathbf{c}}^{t}_{i}\}_{i=1}^{N(t)}$), for object proposals $\{o^{t}_{i}\}_{i=1}^{N(t)}$ in the observed frames. We then compute the matrix $\mathbf{V}_{i}$ by stacking visual features $\{\mathbf{v}_{i}^{t}\}_{t=1}^{T}$. We employ a transformer encoder [vaswani_et_al_2017] to aggregate temporal information across extracted visual features from all frames. The input for our encoder consists of sequences $\mathbf{V}_{i}$, which serve simultaneously as queries $\mathbb{Q}$, keys $\mathbb{K}$ and values $\mathbb{V}$. Formally, the transformation process at the $n$-th encoder layer is given by
@@ 5.2 Spatial Context Processing Unit
To learn spatial context-aware relationship representations. We first construct relationship representations between interacting objects in a scene and further process them by passing through an encoder. Specifically, following [cong_et_al_sttran_2021], we construct relationship representation $\mathbf{z}_{ij}^{t}$ by concatenating visual and semantic features of the objects as follows:
@@
Here, $\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{W}_{3}$ represent the learnable weights of linear layers. $\mathbf{U}_{ij}^{t}$ represent processed feature maps of the union box computed by RoIAlign [ren2016faster]. $\mathbf{S}_{i}^{t}$ and $\mathbf{S}_{j}^{t}$ correspond to the semantic embedding vectors of the object categories. Subsequently, we employ a transformer encoder to integrate spatial contextual information. Here, for each observed frame $I^{t}$, we construct $\mathbf{Z}^{t}$ by stacking all relationship features $\{\mathbf{z}_{ij}^{t}\}_{ij}$ corresponding to objects observed in a frame. We then feed it as input for a transformer encoder which operates on queries $\mathbb{Q}$, keys $\mathbb{K}$, and values $\mathbb{V}$ that are derived from the same source, the preceding layer’s output or the initial relationship features. Formally, the transformation process at the $n$-th encoder layer is given by:
${[\mathbf{Z}^{t}}]^{(n)}=\operatorname{SpatialEncoder}\left(\mathbb{Q}=\mathbb{K}=\mathbb{V}=[{\mathbf{Z}^{t}}]^{(n-1)}\right)$ (5)
@@ 5.3 Latent Dynamics Processing Unit
To understand spatio-temporal dynamics of the evolution of relationships between interacting objects, departing from traditional approaches that architectural variants of transformers [girdhar2021anticipative, thickstun2023anticipatory], we take a continuous time approach and learn a governing differential equation in latent space of relationship representations. Our LDPU contains two components: (1) Temporal Context Encoder and (2) Latent Dynamics Model.
Temporal Context Encoder. To learn spatio-temporal context-aware relationship representations, we further process the output of SCPU by passing it through a transformer encoder that fine-tunes the representation. Specifically, we compute the matrix $\mathbf{Z}_{ij}$ by stacking the relationship representations $\{\mathbf{z}_{ij}^{t}\}_{t=1}^{T}$. We feed $\mathbf{Z}_{ij}$ through an encoder that aggregates temporal information across all observed relationship representations refer Eq. 6. Here, the input $\mathbf{Z}_{ij}$, serves simultaneously as queries $\mathbb{Q}$, keys $\mathbb{K}$, and values $\mathbb{V}$.
@@
Latent Dynamics Models. We begin by abstracting the complexities introduced by external factors that add uncertainty to the evolution of relationships between interacting objects and construct a model based on the premise that the core dynamics driving the changes in these relationships are governed by a non-linear deterministic process. Specifically, we leverage the expressive nature of NeuralODEs [Chen_2018] in learning time-invariant vector fields from data that approximate the underlying non-linear deterministic process and propose SceneSayerODE. We cast the problem of anticipating relationship representations of interacting objects as an instance of the ODE-IVP, where the initial condition is given by the relationship representation of the last observed interaction of the object pair $\left(o_{i}^{t},o_{j}^{t}\right)$. The evolution of this relationship representation over time is then mathematically described as:
@@
In video data that is captured from a single viewpoint, we lose information due to issues such as blurry imagery and occlusions, leading to uncertain interpretations of the scenes. Therefore, it is crucial to integrate uncertainty into the modelling frameworks to represent these stochastic dynamics accurately. We assume the presence of a non-linear stochastic differential equation that governs this evolution and propose SceneSayerSDE which uses NeuralSDEs [kidger2021efficient] to learn it from data. Specifically, we formulate the problem of anticipating future interactions as an SDE-IVP. Here, the initial conditions are set using the representations of the last observed representation object pair, thus facilitating a data-driven learning process that accounts for the inherent uncertainties. The evolution of relationship representation over time is then described as:
@@ 5.4 Decoders
  • Predicate Classification Head. We employ two predicate classification heads (two-layer multi-layer perceptron) in our approach; we use one head for classifying relationship representations between objects of the observed scenes and the other head for classifying anticipated relationship representations of the future scenes.
  • Bounding Box Regression Head. We employed a regression head (two-layer multi-layer perceptron) that takes the anticipated relationship representations as input and outputs bounding box information corresponding to the interacting objects [^4] We observed that decoding bounding boxes of only actors also produced compelling results..
  • Object Classification Head. We decode the output of the object encoder using a two-layer multi-layer perception to generate object category distributions.
@@
Our training objective is constructed as a combination of loss functions implemented on two types of representations (corresponding to objects and relationships): (a) those derived from observed frames, and (b) those anticipated based on these observed frames.
(a) Observed Representations. We apply the following loss functions over representations derived from observed frames: (I) Object Classification Loss on the object representations and (II) Predicate Classification Loss on the relationship representations.
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t},\quad\mathcal{L}_{i}^{t}=-\sum_{n=1}^{|\mathcal{C}|}y_{i,n}^{t}\log(\hat{\mathbf{c}}_{i,n}^{t})}_{\text{Object Classification Loss (I)}};\quad\underbrace{\mathcal{L}_{\text{gen}}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{\text{gen}}^{t},\quad\mathcal{L}_{\text{gen}}^{t}=\sum_{ij}\mathcal{L}_{p^{t}_{ij}}}_{\text{Predicate Classification Loss (II)}}$
(I) Object Classification Loss ($\mathcal{L}_{i}$). We evaluate object classification performance by applying cross-entropy loss to the outputs from the ORPU as detailed above where $y_{i}$ represents the target for object $i$, and $\hat{\mathbf{c}}_{i}^{t}$ is the distribution over object categories.
(II) Predicate Classification Loss ($\mathcal{L}_{\text{gen}}$). focuses on classifying the relationship representations between pairs of objects $\left(o_{i}^{t},o_{j}^{t}\right)$ across all frames ($t\in[1,\bar{T}]$) as detailed above. Here, $\mathcal{L}_{p^{t}_{ij}}$ represents multi-label margin loss and is computed as follows:
$\mathcal{L}_{p^{t}_{ij}}=\sum_{u\in\mathcal{P}^{+}}\sum_{v\in\mathcal{P}^{-}}\max(0,1-\hat{\mathbf{p}}^{t}_{ij}[v]+\hat{\mathbf{p}}^{t}_{ij}[u])$ (9)
(b) Anticipated Representations. We apply the following loss functions over anticipated relationship representations for each observation window: (III) Predicate Classification Loss, (IV) Bounding Box Regression Loss, and (V) Reconstruction Loss
(III) Predicate Classification Loss ($\mathcal{L}_{\text{ant}}^{(1:T)}$). focuses on classifying the anticipated relationship representations between observed pairs of objects $\left(o_{i}^{t},o_{j}^{t}\right)$ in each window. We utilize the frames observed in each window ($t\in[1,T]$) to anticipate the relationship representations in subsequent future frames ($t\in[T+1,\min(T+H,\bar{T})]$) where $T,\bar{T}$ and $H$ denote the number of frames observed in the window, the total number of frames in the video and the length of the anticipation horizon respectively.
$\mathcal{L}_{\text{ant}}^{(1:T)}=\sum_{t=T+1}^{\min(T+H,\bar{T})}\mathcal{L}_{\text{ant}}^{t},\quad\mathcal{L}_{\text{ant}}^{t}=\sum_{ij}\mathcal{L}_{p^{t}_{ij}}$ (10)
(IV) Bounding Box Regression Loss ($\mathcal{L}_{\text{boxes}}^{(1:T)}$). We input anticipated relationship representations into a dedicated linear [^5] In AG, dedicated refers to the allocation of decoders: one for subject and one for objects. layer to estimate the bounding boxes for objects. Inspired by the YOLO models, we compute the difference between the predicted and the ground truth bounding boxes using the Smoothed L1 loss as follows:
$\mathcal{L}_{\text{boxes}}^{(1:T)}=\sum_{t=T+1}^{\min(T+H,\bar{T})}\mathcal{L}_{\text{boxes}}^{t},\quad\mathcal{L}_{\text{boxes}}^{t}=\sum_{k\in\text{boxes}}\text{L}_{\text{smooth}}(b_{k}^{t}-\hat{b}_{k}^{t})$ (11)
(V) Reconstruction Loss ($\mathcal{L}_{\text{recon}}^{(1:T)}$). We apply a Smoothed L1 loss to ensure that the anticipated representations mirror the output from the temporal encoder in LDPU.
$\mathcal{L}_{\text{recon}}^{(1:T)}=\sum_{t=T+1}^{\min(T+H,\bar{T})}\mathcal{L}_{\text{recon}}^{t},\quad\mathcal{L}_{\text{recon}}^{t}=\frac{1}{N(t)\times N(t)}\sum_{ij}^{(N(t)\times N(t))}\text{L}_{\text{smooth}}(\mathbf{z}_{ij}^{t}-\hat{\mathbf{z}}_{ij}^{t})$ (12)
@@
$\mathcal{L}=\underbrace{\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathcal{L}_{\text{gen}}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)}_{\text{Loss Over Observed Representations}}+\underbrace{\sum_{T=3}^{\bar{T}-1}\left(\lambda_{3}\mathcal{L}_{\text{ant}}^{(1:T)}+\lambda_{4}\mathcal{L}_{\text{boxes}}^{(1:T)}+\lambda_{5}\mathcal{L}_{\text{recon}}^{(1:T)}\right)}_{\text{Loss Over Anticipated Representations}}$ (13)
@@ 6 Experiments
Dataset. We apply the proposed method to anticipate future interactions on the Action Genome [Ji_2019]. We pre-process the data and filter out videos with less than 3 annotated frames. Thus, we obtained 11.4K videos in total; we adhered to the train and test split provided by the dataset. The dataset encompasses 35 object classes and 25 relationship classes. These 25 relationship classes are grouped into three categories, namely: (1) Attention Relations comprise relationship classes which primarily describe attention of the subject towards the object, (2) Spatial Relations comprise relationship classes that describe the spatial relationship between two objects, and (3) Contacting Relations comprises relationship classes that indicate different ways the object is contacted.
@@
Remark. This dataset primarily includes videos featuring a single actor engaging with objects [^6] Adopting an actor-centric approach aligns with the literature on VidSGG. While SceneSayer can anticipate relationship representations between any two objects in a scene, we specifically tailor it to address the dynamics between the interacting actor (subject) and the objects. in various real-world environments. In the context of the action genome, a subject-object pair can demonstrate multiple spatial and contacting relationships.
@@
Evaluation Metric. We evaluate our models using the standard Recall@K and meanRecall@K metrics, where $K$ is set to values within the set {10, 20, 50}. Recall@K metric assists in assessing the ability of our model to anticipate the relationships between observed objects in future frames. The long-tailed distribution of relationships in the training set [nag_et_al_tempura_2023] can generate biased scene graphs. While the performance on more common relationships can dominate the Recall@K metrics, the mean recall metric introduced in [chen2019knowledgeembedded] is a more balanced metric that scores the model’s generalisation to all predicate classes. In the following sections, we will assess the performance of models by modifying the initial fraction of the video, denoted as $\mathcal{F}$ provided as input. We set $\mathcal{F}$ to 0.3, 0.5, 0.7, and 0.9 to facilitate a comprehensive understanding of the model’s proficiency in short-term and long-term relationship anticipation.
Settings. We define three settings to evaluate models for the SGA task, each varying by the amount of scene information provided as input:
  • Action Genome Scenes (AGS): In AGS, the model’s input is limited to raw frames of the video. Thus, with minimal context and without any additional information we intend to challenge the model’s ability to interpret and anticipate the scenes.
  • Partially Grounded Action Genome Scenes (PGAGS): In this intermediate setting, along with raw frames of the video we additionally input the model with precise bounding box information of active interacting objects observed in the scene.
@@
  • Grounded Action Genome Scenes (GAGS): In GAGS, we provide the most comprehensive level of scene information. Specifically, the model takes precise bounding box information and the categories of the active objects observed in the interaction as input. In video data, observed frames are fraught with challenges such as object occlusions and blurry imagery, which complicate object detection and interpretation of interaction. Thus, in a setting where we provide complete information regarding a scene as input and reduce the noise induced by these challenges, we aim to evaluate the model’s ability to understand spatio-temporal dynamics.
Figure 3: Overview of Baselines. In our proposed Variant 1 (shown to the left), we input relationship representations to an anticipatory transformer to generate relationship representations for future frames auto-regressively. A predicate classification network (MLP) is then employed to decode these anticipated relationship representations. Meanwhile, in Variant 2 (shown to the right), we enhance relationship representations by passing them through a temporal encoder. These representations are fed to an auto-regressive anticipatory transformer to anticipate future relationship representations. Here, we employ two predicate classification heads (MLPs): one decodes the observed relationship representations, and the other decodes the anticipated relationship representations. In both variants proposed, the auto-regressive anticipatory transformer acts as the generative model, predicting the evolution of relationships between the interacting objects.